OpenStudy (anonymous):

expand.. please tell me why there is 89 people online and not 1 has helped me on this problem... getting fiesty :/

OpenStudy (anonymous):

\[\ln \sqrt[3]{x^2y}\]

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

could you enlarge the text?

OpenStudy (anonymous):

how do i do that

OpenStudy (lgbasallote):

\[\huge \ln \sqrt[3]{x^2y}?\]

OpenStudy (anonymous):

\[\huge \ln \sqrt[3]{x^2y} \]

OpenStudy (lgbasallote):

sorry i gotta go...lemme call on @ParthKohli for you

Parth (parthkohli):

\(\color{Black}{\Rightarrow \ln \sqrt[3]{x^2y} = \ln (x^2y)\large ^{1 \over 3} }\) right?

OpenStudy (anonymous):

okay thanks @igbasallote

Parth (parthkohli):

\(\color{Black}{\Rightarrow \Large {1 \over 3}\ln x^2y }\) \(\color{Black}{\Rightarrow \Large {1 \over3} \ln (xy^{1 \over 2})^2 }\) \(\color{Black}{\Rightarrow {2 \over 3}\Large \ln x\sqrt y }\)

OpenStudy (anonymous):

how did you get that?

Parth (parthkohli):

Do you know the properties of logarithms?

OpenStudy (anonymous):

everyone has asked that? i dont think i know them...

Parth (parthkohli):

Okay, her'es a vital one: \(\color{Black}{\Rightarrow \ln x^y = y \ln x}\)

Parth (parthkohli):

Well you could just simplify it as: \(\color{Black}{\Rightarrow \Large {1 \over 3} \ln x^2y }\)

OpenStudy (anonymous):

ohhh okay!

OpenStudy (anonymous):

thank youuuuuu

Parth (parthkohli):

Even I had no idea what I was doing there lol but it's all good.

OpenStudy (anonymous):

\[2lnx + 3lny\]